Two masses are tied together and one is tied to a
Answer #1

Angular velocity is

\omega =100 rev/min=\frac{100}{60}\times2\pi =10.472 rad/s

Centripetal acceleration for m will be

a=r_1\omega^2=0.10\times10.472^2=10.97m/s^2

Equation for mass mass m1 is

T_1-T_2=m_1r_1\omega^2

and for mass m2 is

T_2=m_2r_2\omega^2=2\times0.20\times10.472^2=43.865

By putting value of T2 we get

T_1=T_2+m_1r_1\omega^2=43.865+2\times0.10\times10.472^2=65.8N